OpenStudy (anonymous):

I need help with this algebra 2 question PLEASEEE

OpenStudy (anonymous):

What is the possible discriminant of the graph? a parabola opening up x intercepts near negative 7 and negative 3 –13 0 15 16

OpenStudy (anonymous):

Discriminant is when y = 0 right? I don't really know.

OpenStudy (tkhunny):

"Near" is kind of an odd instruction. But if we play like it is EXACTLY -7 and -3, we might get (x+7)(x+3) = x^2 + 10x + 21 and the discriminant is 100 - 4(1)(21) = 16 Tricky little problem. I've actually never seen one like it. I like this one. It requires some thought and imagination.

OpenStudy (anonymous):

That's the discriminant? I see.

OpenStudy (anonymous):

ohh okay thank you